OpenStudy (anonymous):

alright mates, gonna need some help on this what is the solution of the equation 9^(2x)=6? round answer to the nearest ten thousandth

OpenStudy (phi):

take the log (base 10 is ok) of both sides \[ \log_{10}\left(9^{2x}\right) = \log_{10}(6) \] use the property \[ \log(a^b) = b \log(a) \] to "get at" the 2x you will need a calculator to find a value for x.

OpenStudy (anonymous):

alriiiiight haha, seeing as how i don't have a calculator is there another way to do this?

OpenStudy (phi):

you can type the problem into google. but first you have to solve for x. can you do that ?

OpenStudy (anonymous):

nope :P i did attempt several times before posting my question but i cant math :P

OpenStudy (phi):

match up this \[ \log(a^b) = b \log(a) \] with the left side of your problem \[ \log\left(9^{2x}\right) = \log(6) \] the idea is to rewrite the left side, using that rule

OpenStudy (anonymous):

alright thanks :)

OpenStudy (phi):

can you rewrite the left side, using that rule ?

OpenStudy (anonymous):

i can try xD

OpenStudy (anonymous):

Look at it this way. We're 'shifting' the \(2x\) to the front.